Project

General

Profile

Actions

20150119 FF

Phone

http://mice.iit.edu/phonebridge.html

Failing that:

Backup Phone Number

Agenda

  • Run plan for 21st Jan
  • Report at MICE Video Conference
    • Please produce 5 slides - preparations/issues during run up, successes, failures
    • Rogers will cover PPS, TOF, KL, Ckov (if there is anything to report)
  • Status, by system (as before)

VC agenda

15:00 - 15:20 Spokesperson's Update 20' Speaker: Prof. Kenneth Long (Imperial College London)
15:20 - 15:40 Project Manager's Report 20' Speaker: Roy Preece (STFC Rutherford Appleton Laboratory)
15:40 - 16:00 MOM Report 20' Speaker: Dr. Chris Rogers (STFC)
16:00 - 16:10 Mock Run - Online 10' Speaker: Edward Overton (The University of Sheffield)
16:10 - 16:20 Mock Run - DAQ 10' Speaker: Yordan Ivanov Karadzhov (Universite de Geneve (CH))
16:20 - 16:30 Mock Run - EMR 10' Speaker: Francois Elie Rene Drielsma (Universite de Geneve (CH))
16:30 - 16:40 Mock Run - Tracker 10' Speaker: David Adey (Fermilab)
16:40 - 16:50 Mock Run - Beamline & Datamover 10' Speaker: Henry Nebrensky (Brunel University)
16:50 - 17:00 Mock Run - Beamline Controls/Controls 10' Speaker: Pierrick Hanlet (Illinois Institute of Technology)

Notes

09:00 start
Rogers will be responsible for
  • taking notes
  • keeping things running
    • watching the clock
    • watching the run plan
    • judging when to halt at a problem and when to progress
  • Kicking folks out of MLCR
PPS
  • PPS issue some problem in the trench
  • Not allowed to run PPS anyway
  • Can't open beamstop, can run D1, D2 and target
Target
  • target frame not tested, frame expected to work
  • would like to run a bit before frame is lowered
  • would like to monitor the target temp
  • gate valve should be open, but want to double check - implication for ISIS happiness ACTION: Rogers
Decay solenoid
  • Working as of end of last week
  • Power supply needs rebooting on Tuesday
NC Magnets
  • Q1 complaining about water on Thurs, fluctuations in readout
  • D1, D2 depend on PPS, run in muon beam permit should be okay

Beamstop okay

Proton absorber okay

Diffuser okay

TOF and KL
  • HV control is fixed
  • SL5.7 machine installed and deployed, CAEN will not support for SL6.4
  • Discussion regarding HV power supplies
  • Check TAG bar channels in HV for KL are turned off ACTION Hanlet - Done
Tracker
  • Four cryostats are cold enough
  • Leak issues
  • Fine temp control is awry, can't self regulate properly, feedback loop is possibly too long but debugging equipment is not available (PC dying), Ed/Craig/Paul trying to improve debugging, too cold, need to warm up with heaters, nominally working at the moment, alternatively use fine control but being "difficult".
  • Power supply unit on cryo 3 is busted, Craig to move power supply from 1/2 to 3
  • Noise issue on Cryo 1/2, originally solved by rerouting power and sensor wires, solved for 1/2, expect 3/4 to be okay based on monitoring of the cables
  • Controls okay, running on SL6.4 still purely Leaver software
    • Board no longer hangs on firmware upgrade, dodgy card, means we can only readout one tracker, two cryostats, sent back to CAEN, spare ordered
  • Trigger - Yordan is now happy, should work, looked good on scope, ought to work
  • DAQ - only two functioning VME boards with upgraded firmware, one for C+M, one for readout
  • DATE - Tried to catch Yordan last week, Yordan was not able to integrate the code, he prescribed a major refactor, major refactor failed, David Adey will implement a wrapper to keep tracker code independent of DATE
  • Unpacking still blinged
    • Attempting to write code to single station test unpacking
    • Online monitoring which existed last time has disappeared - Ed Santos standalone has gone
    • Online recon which existed last time has disappeared
    • Melissa is making progress but hitting technical difficulties; very tight timescale
  • LED - not seen in November, not attempted it yet this week, too many other issues
    • Move towards LED, abandon standalone system
  • Craig - do we need to do calibration? David Adey - No it is okay, we should see something regardless, calibration is too hard anyway
  • LM power supply - needs checking voltages, connectors have been pulled out
    • Pierrick proposes monitoring
  • EMR, same as before
  • DAQ, dropping out once in a while, dropping broadcasts, nothing changed with DATE version, possibly problem with network, Ed to sit next to Yordan, but rather complex
Controls
  • Controls beamline IOC is complete, tested, except one small bug
    • Decay Solenoid and Target separate by design
  • Beamline state machine operational but not fully operational, need to run the beamline on Wednesday
  • Beamline run control is refactoring work in progress
  • Detectors, a lot to do with EMR, integrated detector ioc will not be implemented
    • 527 and 127 are operational
    • New PC running the high voltage, SL5.7, PCI card running
  • Tracker uses repurposed MICEACQ22
    • Some discussion about spares etc
  • TestIOCPC code is moved into production on MICEIOCPC1
  • ACTION Rogers to talk to Craig about spares with MOM hat on
MLCR Servers
  • Nothing new has died, broken things have been fixed
  • MICEIOCPC1 RAID died on two disks
Datamover
  • Okay, not tested since network switch was changed, PPD network communications switched
  • ACTION Janusz to check
  • Still moving on manual
  • Henry needs to check that mount points are on MICE RAID machines using automount daemon
  • Data size threshold will stay in for now
  • Fast turnaround reconstruction, work in progress, different path at RAL, one bit missing, writing a test script
    • Fall back is to reconstruct through Brunel/IC
  • NAGIOS monitoring not implemented
Online Recon
  • Using old version, with old unpacking, should work as not touched since last time
  • Need unpacking to work prior to March
  • Durga/Dobbs will install a new version e.g. 0.9.2 on Onrec machines
Online Monitoring
  • Rhys is apparently at RAL but not been seen
Henry - documentation
  • Target check list manual is up to date
  • NC magnet manual is a bit old
  • DS is not documented
  • Beamstop docs are up to date
  • PAbs docs are missing
  • Diffuser docs are missing
  • Movefiles docs are missing

ACTION Rogers: Run conditions spreadsheet should be used to check CDB write

Note data okay flag

Long term actions:

  • Make R76 a viable place for MICE S/w&C pied a terre
  • ToF/KL HV supply issue: needs new system
  • Some old PCs may still be in use; need to purchase replacements;
  • Automation of data moving

Updated by Rogers, Chris over 8 years ago ยท 18 revisions